OpenStudy (anonymous):

What is the sum of the roots of the following equation? 3x^2 − 24 = 6x (A) −4 (B) −2 (C) 2 D) 4

OpenStudy (alexandervonhumboldt2):

ok so 3x^2-6x-24=0 use the quadratic formula http://www.statisticslectures.com/images/quadeq1.gif (6+-sqrt(6^2+4*24*3))/6=(6+-18)/6=-2, 4 -2+4=2

OpenStudy (alexandervonhumboldt2):

\(\Huge\text{Welcome to OpenStudy!!!}\)

OpenStudy (anonymous):

i don't get this part (6+-sqrt(6^2+4*24*3))/6=(6+-18)/6=-2, 4

OpenStudy (anonymous):

it doesn't pop up properly

OpenStudy (tkhunny):

Ponder: \((x-a)(x-b) = x^{2} + (a+b)x + ab\)

OpenStudy (anonymous):

how do i do that?

OpenStudy (tkhunny):

Subtract 6x